OpenStudy (anonymous):

The Ross family recently dinned at a Italian restaurant.Each of the four family members ordered a pasta dish that cost $11.50, and a drink that cost $1.50, and a dessert that cost $2.75. write an expression that could be used to calculate the cost of the Ross' dinner before adding tax and a tip.

OpenStudy (anonymous):

4( pasta + drink + dessert ) 4 ( 15.75) 63

OpenStudy (anonymous):

P (x) = 11.50x + 1.50x + 2.75x P (x) = 15.75x x = number of people in family :D P (x) = 15.75 * 4 = $63